# A Framework For Automated Dissection Along Tissue Boundary

### v3

@@ I Introduction
The main sensing modality in RAS is the cameras and there is considerable research on processing these images. In [10], the creation of point clouds allows for a 3D modeling of the surgical site. These models can then be used to guide the robot’s movements. The ability to generate and process point clouds is especially crucial when dealing with deformable objects, as it provides a dynamic representation that can adapt to changes in real-time. However, the point clouds in [10] are created from plenoptic cameras which are not present in most operating rooms. Furthermore, the addition of a second camera to the surgical setup is not compatible with the current robot-assisted laparoscopy, which relies on endoscopic images. Other studies have employed visual segmentation models to detect and define loose connective tissues, aiding surgeons in visualizing safe dissection planes during procedures like robot-assisted gastrectomy [11]. Nevertheless, the processing is at the pixel level, and there is no segmentation and logical connection to individual tissues.
Compared to auto-suturing, there seems to be less work on auto-dissection, indicating a potential research gap [12]. However, [13] introduces a vision-based cutting control algorithm, emphasizing the role of visual servoing in enhancing the precision of automated dissection. This algorithm allows for the automatic cutting of deformable objects along a predetermined path with a scalpel, utilizing visual feedback from feature points for real-time parameter estimation. In RAS, surgeons never use scalpels due to bleeding and use energy delivery with monopolar, bipolar, or ultrasonic instruments, instead.
@@
In our work, we use the da Vinci surgical robot with the da Vinci Research Kit (dVRK) [14]. Our research focuses on vision-based robot control tailored to tissue dissection [^1] In the surgical literature, dissection refers to the actual separation of tissues; in our work, the term refers to energy delivery to a particular location on a specimen.. Our main contributions are: (a) We introduce a novel comprehensive framework for dissecting the tissue relying solely on endoscopic images, that is what is provided to a surgeon, to control the robot. (b) To support precise robot control, we developed an ArUco-based [15] calibration for dVRK, which is easy to implement and remains feasible when the configuration of the da Vinci system is changed. (c) We augmented the dVRK to control the power supply that delivers energy to the monopolar and bipolar instruments for dissection. This allows us to use additional instruments like the Permanent Cautery Hook (PCH), which is rarely used in research but widely used in cholecystectomy. (d) We generated a custom dataset that categorizes different types of tissues and da Vinci instruments. This dataset was used for fine-tuning Detectron2 [16] and we show that the new model can be used in RAS for segmenting tissues and detecting keypoints of instruments. This allows the system to identify the boundary between different tissues and locate the instrument. (e) We conducted an ex-vivo evaluation of our proposed system using pig liver and chicken specimens.
We do not claim that we have achieved realistic tissue dissection as we do not address tissue flexibility or use a grasper to pull the tissue during the dissection; these are left for future work. However, we show that we can successfully automate a number of fundamental subtasks during tissue dissection, and for the first time demonstrate a robotic system that can autonomously deliver energy along tissue boundaries. Our work is therefore an important step towards autonomous dissections that approach surgeon’s skill.
@@ II-B Custom Calibration
Measure (distance) dVRK normal-$\Leftrightarrow$ ArUco Custom normal-$\Leftrightarrow$ ArUco
@@ II-C Image Segmentation and Keypoint Detection
The core challenge in automating surgical procedures is accurately extracting surgical instrument locations and locations of target tissues from the endoscopic images. To achieve this, we first curated a custom dataset employing the Segment Anything Model (SAM) which was pre-trained with the Segment Anything 1 Billion (SA-1B) [20]. It utilizes prompts defining the segmentation targets in an image, suitable for a broad spectrum of segmentation tasks without requiring further training. Manually selected points were provided as prompts, which are categorized as either positive or negative to indicate whether they belong to the foreground (the object of interest) or the background. The project couldn’t use existing segmentation datasets such as CholecSeg8k  [21] or CholecTriplet2021  [22] because they’re based on in-vivo human liver surgeries, while this study involves ex-vivo porcine models. Differences in anatomy and surgical environment between humans and pigs, as well as in-vivo and ex-vivo settings, affect tissue and instrument appearance. Furthermore, providing a streamlined workflow that easily generalizes to other tissues and procedures is an important contribution of our work.
@@
  • TipRight/TipLeft located for the Fenestrated Bipolar Forceps (FBF) and Large Needle Driver (LND) at the two extremities of the grippers, and for the Permanent Cautery Hook (PCH) at the tip and beginning of the hook.
@@
These key points are on key parts of the tools that have different colors and distinctive edges to maximize their detection. We recorded endoscopic images while performing motions with both arms. To ensure data diversity, the endoscope angle and the position of the ex-vivo material (chicken or pig liver) were changed several times, resulting in a total of 2820 images. Details of the number of instances annotated for segmentation and keypoint detection are described in Table II. The dataset adheres to Microsoft’s COCO format [24], ensuring compatibility and ease of integration.
As shown in the first two rows of Fig. 4, from semi-automatically annotated images for image segmentation, and manual annotations for keypoint detection (Fig. 4(a)) it was then possible to obtain a precise (cf. Table III) model generalizable to different shapes of specimens (pig livers and chicken) and instrument configurations.
The annotated dataset was trained on Detectron2 [16], a state-of-the-art object segmentation and keypoint detection model. The training was based on the pre-trained mask R-CNN model R50-FPN since it had the fastest inference time among different models with similar performance. The learning rate was initialized as 0.05 with a warm-up phase of 1000 iterations. The model was trained for 9000 iterations to ensure comprehensive learning and the learning rate was reduced by $20\%$ every 300 iterations. Additionally, a batch size of 512 for the Region of Interests aids in capturing more object instances. These hyperparameters represent a balanced approach to achieving accurate and efficient image segmentation, taking into account computational constraints and dataset characteristics. Similar hyperparameters have been used for keypoint detection.
@@
A significant component of our methodology is the 3D reconstruction of the surgical scene. Before the reconstruction, we captured multiple images including a $9\times 6$ chessboard in different positions and orientations. The cameras were calibrated with selected low reprojection error images according to the traditional approach in [25] from the MATLAB Stereo Camera Calibration Toolbox and OpenCV [26]. Subsequently, using the intrinsic and extrinsic camera parameters, we applied the modified Semi-global Matching algorithm (SGM) [27] to produce stereo disparity maps from the stereo endoscopic images. The images are passed through a bilateral filter before applying SGM to reduce the noise but maintain the edges as much as possible. The disparity map is then converted to point clouds using the baseline and the focal length of the stereo cameras. We average each 3D tissue boundary and instrument position point within a fixed-size ($10\times 10$) window, thereby identifying points representing infinite distances and eliminating them from the disparity map as outliers.
@@ II-E Control Algorithm
Our approach leverages Detectron2 for processing endoscopic images, differentiating between dual targets like muscle and skin in chicken samples. The adjacent points were found by the traditional method for finding nearest neighbors with K-Dimensional Tree (KD Tree) [28]. Subsequently, the edge point between each set of background tissue (e.g. liver) and the primary tissue (e.g. gallbladder) trajectory point was extracted based on the peak disparity value between the two points. Afterward, the edge points were downsampled based on the farthest-first traversal algorithm [29].
@@ III-B Ex-vivo Evaluation
Trial Mean (mm) Standard deviation (mm)
Liver 1 0.31 0.23
2 0.37 0.27
3 0.29 0.24
4 0.35 0.28
5 0.30 0.20
6 0.33 0.19
Chicken 1 0.28 0.18
2 0.59 0.72
3 0.35 0.25
4 0.41 0.35
5 0.27 0.17
Weighted 0.36 0.34
@@ IV Conclusion
Furthermore, the 3D scene reconstruction, achieved through the Semi-global Matching algorithm (SGM) on stereo endoscopic images, along with our fiducial marker-based calibration has enabled real-time monitoring of the instrument’s pose. This, in tandem with our feedback mechanism and control strategy, ensures the robot’s precise trajectory following during the dissection process.
@@
To resolve the issues, we collected a set of ex-vivo cholecystectomies performed by surgeons on our current da Vinci model [30]. Each trial was carried out with a different specimen and diverse movements of the instruments were observed. The future model will be trained with this augmented dataset, expecting it to be rich enough to produce a real-time algorithm to robustly track the tissue boundary and plan instrument movement. This will allow the system to respond when the gallbladder is grasped and pulled, or if the gallbladder deforms as a result of the energy delivery. Moreover, we will implement the endoscope movement as in [31] to keep the instrument consistently centered. We also plan to initially combine manual control of the grasper with automated boundary tracking and energy delivery, as automating the grasping task is significantly more challenging than dissection. However, our long-term goal is to completely automate the dissection task with performance similar to that of a surgeon.

### v4

@@ I Introduction
The cameras are the main sensing modality in RAS, and considerable research has been done on processing these images. In [10], creating point clouds allows for a 3D modeling of the surgical site. These models can then be used to guide the robot’s movements. The ability to generate and process point clouds are especially crucial when dealing with deformable objects, as it provides a dynamic representation that can adapt to changes in real-time. However, the point clouds in [10] are created from plenoptic cameras, which are not present in most operating rooms. Furthermore, the addition of a second camera to the surgical setup is not compatible with the current robot-assisted laparoscopy, which relies on endoscopic images. Other studies have employed visual segmentation models to detect and define loose connective tissues, aiding surgeons in visualizing safe dissection planes during procedures like robot-assisted gastrectomy [11]. Nevertheless, the processing is at the pixel level, and there is no segmentation and logical connection to individual tissues.
Compared to auto-suturing, there seems to be less work on auto-dissection, indicating a potential research gap [12]. However, [13] introduces a vision-based cutting control algorithm, emphasizing the role of visual servoing in enhancing the precision of automated dissection. This algorithm allows for the automatic cutting of deformable objects along a predetermined path with a scalpel, utilizing visual feedback from feature points for real-time parameter estimation. In RAS, surgeons never use scalpels due to bleeding and use energy delivery with monopolar, bipolar, or ultrasonic instruments instead.
@@
In our work, we use the da Vinci surgical robot with the da Vinci Research Kit (dVRK) [14]. Our research focuses on vision-based robot control tailored to tissue dissection [^1] In the surgical literature, dissection refers to the actual separation of tissues; in our work, the term refers to energy delivery to a particular location on a specimen.. Our main contributions are: (a) We introduce a novel comprehensive framework for dissecting the tissue relying solely on endoscopic images, that is, what is provided to a surgeon to control the robot. (b) To support precise robot control, we developed an ArUco-based [15] calibration for dVRK, which is easy to implement and remains feasible when the configuration of the da Vinci system is changed. (c) We augmented the dVRK to control the power supply that delivers energy to the monopolar and bipolar instruments for dissection. This allows us to use additional instruments like the Permanent Cautery Hook (PCH), which is rarely used in research but widely used in cholecystectomy. (d) We generated a custom dataset that categorizes different types of tissues and da Vinci instruments. This dataset was used for fine-tuning Detectron2 [16], and we show that the new model can be used in RAS for segmenting tissues and detecting keypoints of instruments. This allows the system to identify the boundary between different tissues and locate the instrument. (e) We conducted an ex-vivo evaluation of our proposed system using pig liver and chicken specimens.
We do not claim that we have achieved realistic tissue dissection as we do not address tissue flexibility or use a grasper to pull the tissue during the dissection; these are left for future work. However, we show that we can successfully automate a number of fundamental subtasks during tissue dissection and, for the first time, demonstrate a robotic system that can autonomously deliver energy along tissue boundaries. Our work is, therefore, an important step towards autonomous dissections that approach surgeon’s skill.
@@ II-B Custom Calibration
Measure (distance) dVRK $\Leftrightarrow$ ArUco Custom $\Leftrightarrow$ ArUco
@@ II-C Image Segmentation and Keypoint Detection
The core challenge in automating surgical procedures is accurately extracting surgical instrument locations and locations of target tissues from the endoscopic images. To achieve this, we first curated a custom dataset employing the Segment Anything Model (SAM), which was pre-trained with the Segment Anything 1 Billion (SA-1B) [20]. It utilizes prompts to define the segmentation targets in an image, which is suitable for a broad spectrum of segmentation tasks without requiring further training. Manually selected points were provided as prompts, which are categorized as either positive or negative to indicate whether they belong to the foreground (the object of interest) or the background. The project couldn’t use existing segmentation datasets such as CholecSeg8k  [21], or CholecTriplet2021  [22] because they’re based on in-vivo human liver surgeries, while this study involves ex-vivo porcine models. Differences in anatomy and surgical environment between humans and pigs, as well as in-vivo and ex-vivo settings, affect tissue and instrument appearance. Furthermore, providing a streamlined workflow that easily generalizes to other tissues and procedures is an important contribution to our work.
@@
  • TipRight/TipLeft located for the Fenestrated Bipolar Forceps (FBF) and Large Needle Driver (LND) at the two extremities of the grippers and for the Permanent Cautery Hook (PCH) at the tip and beginning of the hook.
@@
These keypoints are on key parts of the tools with different colors and distinctive edges to maximize their detection. We recorded endoscopic images while performing motions with both arms. To ensure data diversity, the endoscope angle and the position of the ex-vivo material (chicken or pig liver) were changed several times, resulting in a total of 2820 images. Details of the number of instances annotated for segmentation and keypoint detection are described in Table II. The dataset adheres to Microsoft’s COCO format [24], ensuring compatibility and ease of integration.
As shown in the first two rows of Fig. 4, from semi-automatically annotated images for image segmentation and manual annotations for keypoint detection (Fig. 5(a)) it was then possible to obtain a precise (cf. Table III) model generalizable to different shapes of specimens (pig livers and chicken) and instrument configurations.
The annotated dataset was trained on Detectron2 [16], a state-of-the-art object segmentation and keypoint detection model. The training was based on the pre-trained mask R-CNN model R50-FPN since it had the fastest inference time among different models with similar performance. The learning rate was initialized as 0.05 with a warm-up phase of 1000 iterations. The model was trained for 9000 iterations to ensure comprehensive learning, and the learning rate was reduced by $20\%$ every 300 iterations. Additionally, a batch size of 512 for the Region of Interests aids in capturing more object instances. These hyperparameters represent a balanced approach to achieving accurate and efficient image segmentation, taking into account computational constraints and dataset characteristics. Similar hyperparameters have been used for keypoint detection.
@@
A significant component of our methodology is the 3D reconstruction of the surgical scene. Before the reconstruction, we captured multiple images, including a $9\times 6$ chessboard in different positions and orientations. The cameras were calibrated with selected low reprojection error images according to the traditional approach in [25] from the MATLAB Stereo Camera Calibration Toolbox and OpenCV [26]. Subsequently, using the intrinsic and extrinsic camera parameters, we applied the modified Semi-global Matching algorithm (SGM) [27] to produce stereo disparity maps from the stereo endoscopic images. The images are passed through a bilateral filter before applying SGM to reduce the noise but maintain the edges as much as possible. The disparity map is then converted to point clouds using the baseline and the focal length of the stereo cameras. We average each 3D tissue boundary and instrument position point within a fixed-size ($10\times 10$) window, thereby identifying points representing infinite distances and eliminating them from the disparity map as outliers.
@@ II-E Control Algorithm
Our approach leverages Detectron2 for processing endoscopic images, differentiating between dual targets like muscle and skin in chicken samples. The adjacent points were found by the traditional method for finding nearest neighbors with K-Dimensional Tree (KD Tree) [28]. Subsequently, the edge point between each set of background tissue (e.g., liver) and the primary tissue (e.g., gallbladder) trajectory point was extracted based on the peak disparity value between the two points. Afterward, the edge points were downsampled based on the farthest-first traversal algorithm [29].
@@ III-B Ex-vivo Evaluation
Trial Mean (mm) Std. dev. (mm) Duration (s)
Liver 1 0.31 0.23 128
2 0.37 0.27 138
3 0.29 0.24 102
4 0.35 0.28 113
5 0.30 0.20 117
6 0.33 0.19 135
Chicken 1 0.28 0.18 75
2 0.59 0.72 108
3 0.35 0.25 98
4 0.41 0.35 153
5 0.27 0.17 130
Weighted 0.36 0.34 121.6
@@ IV Conclusion
Furthermore, the 3D scene reconstruction, achieved through the Semi-global Matching algorithm (SGM) on stereo endoscopic images, along with our fiducial marker-based calibration, has enabled real-time monitoring of the instrument’s pose. This, in tandem with our feedback mechanism and control strategy, ensures the robot’s precise trajectory following during the dissection process.
@@
To resolve the issues, we collected a set of ex-vivo cholecystectomies performed by surgeons on our current da Vinci model [30]. Each trial was carried out with a different specimen, and diverse movements of the instruments were observed. The future model will be trained with this augmented dataset, expecting it to be rich enough to produce a real-time algorithm to robustly track the tissue boundary and plan instrument movement. This will allow the system to respond when the gallbladder is grasped and pulled or if the gallbladder deforms as a result of the energy delivery. Moreover, we will implement the endoscope movement as in [31] to keep the instrument consistently centered. We also plan to initially combine manual control of the grasper with automated boundary tracking and energy delivery, as automating the grasping task is significantly more challenging than dissection. However, our long-term goal is to completely automate the dissection task with performance similar to that of a surgeon.